OpenStudy (bimbim2697):

An air gun fires a pellet with a mass of 0.0010 slugs. The slug is fired from the gun, and after which possesses of 720 lb ft of kinetic energy. Find its velocity, in ft/s.

OpenStudy (harman.singh):

Here is the formula for Kinetic Energy: \[KE=\frac{ 1 }{ 2 } \times m \times v^2\] - Solve for velocity(v)

OpenStudy (bimbim2697):

Harman I am completely lost on this one LOL

OpenStudy (harman.singh):

That's completely okay. Here's more information: - KE refers to the Kinetic Energy given to you in the question (720 lb ft) - m is the mass (0.0010 slugs) - v is the velocity that you are trying to find Going back to the equation.. \[KE=\frac{ 1 }{ 2 } mv^2\] \[720=\frac{ 1 }{ 2 } \times 0.0010 \times v^2\] From here on, you can solve for 'v'.

OpenStudy (bimbim2697):

1200

OpenStudy (harman.singh):

Yes, that's correct :)

OpenStudy (bimbim2697):

Thanks so much!

OpenStudy (harman.singh):

You're welcome!

OpenStudy (bimbim2697):

A sandbag is dropped from a hot air balloon at a height of 370 m above the ground. Find its velocity, in m/s, just before it hits the ground. Disregard air resistance.

OpenStudy (bimbim2697):

85 square of 2(9.8)(370)

OpenStudy (harman.singh):

v^2=u^2+2ax v=sqroot of (2*9.8*370( v=85 m/s Yes, your answer is correct!